# Active pattern formation emergent from single-species nonreciprocity

### v1

@@ Abstract
Nonreciprocal forces violating Newton’s third law are common in a plethora of nonequilibrium situations ranging from predator-prey systems to the swarming of birds and effective colloidal interactions under flow. While many recent studies have focused on two species with nonreciprocal coupling, much less is examined for the basic single-component system breaking the actio and reactio equality of force within the same species. Here, we systematically derive the fundamental field theory of single-species nonreciprocal interactions from microscopic dynamics, leading to a generic framework termed Active Model N (N denoting nonreciprocity). We explore the rich dynamics of pattern formation in this intrinsic nonreciprocal system and the emergence of self-traveling states with persistent variation and flowing of active branched patterns. One particular new characteristic pattern is an interwoven self-knitting “yarn” structure with a unique feature of simultaneous development of micro- and bulk phase separations. The growth dynamics of a “ball-of-wool” active droplet towards these self-knitted yarn or branched states exhibits a crossover between different scaling behaviors. The mechanism underlying this distinct class of active phase separation is attributed to the interplay between force nonreciprocity and competition. Our predictions can be applied to various biological and artificial active matter systems controlled by single-species nonreciprocity.
@@
Newton’s famous third law states that the force ${\mathbf{F}}_{1\rightarrow 2}$ which a body 1 exerts on another body 2 is, up to a sign, the same as the force $\mathbf{F}_{2\rightarrow 1}$ which body 2 exerts on body 1 such that there is equality of actio and reactio, i.e., $\mathbf{F}_{1\rightarrow 2}=-\mathbf{F}_{2\rightarrow 1}$. This law of reciprocity holds for all fundamental interactions and also for effective interactions in equilibrium situations. In nonequilibrium, however, reciprocity is typically broken such that $\mathbf{F}_{1\rightarrow 2}+\mathbf{F}_{2\rightarrow 1}\not=\mathbf{0}$ which implies that the center-of-mass of a two-body system is not force-free but experiences a driving force leading to spontaneous “active” motion of a pair.
@@
There are two fundamentally different typical examples of nonreciprocity: either the two bodies are of different species (such as a predator trying to catch an escaping prey) or they are of the same type (such as two birds flying together where only the following bird can see the leading one). In the former case of two-species nonreciprocity typically only particles of different species couple nonreciprocally, while the latter case of single-species nonreciprocity concerns intrinsic nonreciprocal interactions within the same species. There are many examples for single-species nonreciprocity both for living systems such as birds, fish, bacteria [1], or starfish embryos [2], and artificial robotic and colloidal systems with vision-cone-like interactions [3]. Theory has considered various aspects of single-species nonreciprocity stemming from vision-cone effects [4, 5, 6, 7] or transverse interactions [1, 2, 8]. Systems governed by nonreciprocal interactions have shown potential technological importance, as demonstrated in various contexts ranging from robotic metamaterials [9, 10] to drug design [11].
@@
In the field-theoretical description of active matter, most of the recent works on pattern formation induced by nonreciprocal interactions treat coupling-nonreciprocity of a binary system while ignoring nonreciprocity between the same species [12, 13, 14, 15, 16, 17]. These studies have provided crucial insights into the dynamics of multi-component systems [18, 19, 20] and a wealth of novel dynamical behavior has been discovered including binary active crystals [21, 22], symmetry-broken nonreciprocal phase transitions [16], as well as traveling bands [14] and transverse ripples [15].
@@
Here we investigate the intrinsic case of active pattern formation induced by the original form of force nonreciprocity among single species. We start from the explicit form for the nonreciprocal interaction forces ${\mathbf{F}}_{2\rightarrow 1}$ and systematically derive a fundamental underlying continuum field theory based on microscopic dynamics, establishing a new model which we term Active Model N (where N denotes the novel aspect of nonreciprocity). A crucial advantage of this type of continuum field theories [23, 24, 25], in addition to the fact that they can reveal the complex mechanisms involved, is the accessibility to much larger spatial and temporal scales for a greater amount of particles as compared to discrete particle-based simulations. Unlike previous active field theories such as Active Model B [26] and its extensions [27, 28, 29], this model provides a generic framework to study the general class of single-species force nonreciprocity. It is a unifying field theory building an inherent connection between nonreciprocal interaction forces and self-propelled active matter, i.e., matter becoming active by nonreciprocity. We can then predict new types of self-traveling chiral patterns for which both spatial mirror symmetry and time-reversal symmetry are spontaneously broken. In particular, we discover a unique interwoven pattern reminiscent of self-knitting threads that we term “active yarn”, which self-organizes through simultaneous evolution of micro- and bulk phase separations over multiple length and time scales. This new class of active phase separation is induced by an intrinsic mechanism showing as the interplay between interparticle force nonreciprocity and the force competition that is beyond the traditional route, leading to the simultaneous occurrence of two distinct classes of collective behavior incorporating phase-separation-induced assembly and orientationally aligning, persistent collective self-migration. Notably, the size of the phase-separating yarn region is controllable (e.g., via the vision-cone angle). This mechanism has the potential for interesting applications in the realization of programmable matter, particularly in the form of controlled phase separation.
@@
Fig. 1: Schematic of nonreciprocal single-species vision-cone model. An agent 1 with orientation $\mathbf{\hat{u}_{1}}$ sees and interacts with other agents within a vision cone of full opening angle $2\vartheta$. Agent 2 is included in the perception zone of agent 1, such that there is a resulting nonreciprocal force acting on agent 1 which is directed along $\mathbf{r}=\mathbf{r}_{2}-\mathbf{r}_{1}$. There is also a torque turning its orientation to be aligned with the orientation $\mathbf{\hat{u}_{2}}$ of the second agent. In this example, agent 2 with its own vision cone does not see agent 1 and thus there is no interaction between them that influences the motion of agent 2. Likewise, agent 3 is not visible for agent 1 such that there is no mutual interaction between agents 1 and 3 that influences the motion of either of them.
From microscopic to continuum field model
@@
We start from a microscopic picture of particles or agents interacting through vision cones, as illustrated in Fig. 1. Consider a general form of pairwise nonreciprocal force between single-component particles with orientational degree of freedom on a plane, i.e.,
$\mathbf{F}_{2\rightarrow 1}=\hat{\mathbf{r}}f\left(|\mathbf{r}_{2}-\mathbf{r}_{1}|,\hat{\mathbf{u}}_{1}\cdot\hat{\mathbf{r}},\hat{\mathbf{u}}_{2}\cdot\hat{\mathbf{r}}\right),$ (1)
@@
$\mathbf{F}_{2\rightarrow 1}\neq-\mathbf{F}_{1\rightarrow 2}=-(-\hat{\mathbf{r}})f\left(r,-\hat{\mathbf{u}}_{2}\cdot\hat{\mathbf{r}},-\hat{\mathbf{u}}_{1}\cdot\hat{\mathbf{r}}\right),$ (2)
@@
$f(r,\cos\varphi_{1},\cos\varphi_{2})=h(r)\Theta\left(-\cos\vartheta+\cos\varphi_{1}\right),$ (4)
@@
$M(r,\hat{\mathbf{u}}_{1},\hat{\mathbf{u}}_{2})=\tau(r)\Delta\varphi_{21}\Theta\left(-\cos\vartheta+\cos\varphi_{1}\right),$ (5)
@@
To derive Active Model N, we start from the microscopic Langevin equations containing the nonreciprocal forces and torques. The dynamics of the full one-particle density field $\varrho(\mathbf{r},\hat{\mathbf{u}},t)$ is then described by the stochastically equivalent Smoluchowski equation. By a systematic expansion in terms of a particle density field $\rho(\mathbf{r},t)$ and a polarization density field $\mathbf{P}(\mathbf{r},t)$, the coarse-grained continuum field equations [24, 30] are obtained (see Methods and Supplementary Information). We then reach a new continuum density-field theory, i.e., Active Model N,

### v2

@@ Abstract
Nonreciprocal interactions violating Newton’s third law are common in a plethora of nonequilibrium situations ranging from predator-prey systems to the swarming of birds and effective colloidal interactions under flow. While many recent studies have focused on two species with nonreciprocal coupling, less is examined for the basic single-component system breaking the actio and reactio equality of force within the same species. Here, we systematically derive a field theory for the case of single-species nonreciprocal interactions from the microscopic particle dynamics, leading to a generic continuum model termed Active Model N (N denoting nonreciprocity). We explore the rich dynamics of pattern formation in this nonreciprocal system and the emergence of self-traveling states with persistent variation and flowing of active branched patterns. One particular new characteristic pattern is an interwoven self-knitting “yarn” structure with a unique feature of simultaneous development of micro- and bulk phase separations. The growth dynamics of a “ball-of-wool” active droplet towards these self-knitted yarn or branched states exhibits a crossover between different scaling behaviors. The mechanism underlying this distinct class of active phase separation is attributed to the interplay between nonreciprocity and competition of interparticle forces. Our predictions can be applied to various biological and artificial active matter systems controlled by single-species nonreciprocity.
I Introduction
@@
Newton’s famous third law states that the force ${\mathbf{F}}_{1\rightarrow 2}$ which a body 1 exerts on another body 2 is, up to a sign, the same as the force $\mathbf{F}_{2\rightarrow 1}$ which body 2 exerts on body 1 such that there is equality of actio and reactio, i.e., $\mathbf{F}_{1\rightarrow 2}=-\mathbf{F}_{2\rightarrow 1}$. This law of reciprocity holds for all fundamental interactions and also for effective interactions in equilibrium situations. In systems out of equilibrium, however, reciprocity is typically broken such that $\mathbf{F}_{1\rightarrow 2}+\mathbf{F}_{2\rightarrow 1}\not=\mathbf{0}$ which implies that the center-of-mass of a two-body system is not force-free but experiences a driving force leading to spontaneous “active” motion of a pair.
@@
There are two fundamentally distinct typical examples of nonreciprocity: either the two bodies are of different species (such as a predator trying to catch an escaping prey) or they are of the same type (such as two birds flying together where only the following bird can see the leading one). In the former case of two-species nonreciprocity, typically only particles of different species couple nonreciprocally, while the latter case of single-species nonreciprocity concerns intrinsic nonreciprocal interactions within the same species. There are many examples of single-species nonreciprocity, both of living systems such as birds, fish, bacteria [1], or starfish embryos [2], and artificial robotic and colloidal systems with vision-cone-like interactions [3]. Various aspects of single-species nonreciprocity stemming from vision-cone effects [4, 5, 6, 7, 8, 9, 10], transverse interactions [1, 2, 11] or the general cases of parity-symmetric and parity-antisymmetric forces [12] have been considered in theories. Systems governed by nonreciprocal interactions have shown potential technological importance, as demonstrated in various contexts ranging from robotic metamaterials [13, 14] to drug design [15].
@@
In the field-theoretical description of active matter, most of the recent works on pattern formation induced by nonreciprocal interactions treat coupling-nonreciprocity of a binary system while ignoring nonreciprocity between the same species [16, 17, 18, 19, 20, 21]. These studies have provided crucial insights into the dynamics of multicomponent systems [22, 23, 24, 25] and a wealth of novel dynamical behavior has been discovered, including symmetry-broken nonreciprocal phase transitions [20], traveling waves with interface undulations and spatiotemporal chaos [25], as well as traveling bands [18] and transverse ripples [19]. The case of single-species vision-cone nonreciprocity was examined from a field-theoretical perspective in Ref. [4] which, like most studies employing particle-based simulations [4, 5, 6, 7, 8], considered self-propelled particles whose self-propulsion is modified by a vision-cone mechanism.
@@
Here we investigate the rich behavior of pattern formation in a single-species system that becomes active not via self-propulsion of individual particles, but via force nonreciprocity. We start from the microscopic equations of motion for particles interacting via nonreciprocal interaction forces ${\mathbf{F}}_{2\rightarrow 1}$ and systematically derive a fundamental underlying continuum field theory based on microscopic dynamics, establishing a new model which we term Active Model N (where N denotes the aspect of nonreciprocity). A crucial advantage of this type of continuum field theories [26, 27, 28, 29], in addition to the fact that they can reveal the complex mechanisms involved, is the accessibility of much larger spatial and temporal scales for a greater amount of particles compared to discrete particle-based simulations. Unlike previous active field theories such as Active Model B [30] and its extensions [31, 32, 33] as well as active phase field crystal models [34, 35, 36, 37, 38], this model is based on a generic framework to study single-species force nonreciprocity. Through the study of vision-cone-like interactions, it builds an inherent connection between nonreciprocal interaction forces and self-propelled active matter. We predict new types of self-traveling chiral patterns for which both spatial mirror symmetry and time-reversal symmetry are spontaneously broken. In particular, we discover a unique interwoven pattern reminiscent of self-knitting threads that we term “active yarn”, which self-organizes through simultaneous evolution of micro- and bulk phase separations over multiple length and time scales. This new class of active phase separation is induced by an intrinsic mechanism showing as the interplay between interparticle force nonreciprocity and the strength and range of interactions, leading to the simultaneous occurrence of two distinct classes of collective behavior incorporating phase-separation-induced assembly and orientationally aligning, persistent collective self-migration. Notably, the size of the phase-separating yarn region is controllable via, e.g., the vision-cone angle. This mechanism has the potential for interesting applications in the realization of programmable matter, particularly in the form of controlled phase separation.
@@
Figure 1: Schematic of nonreciprocal single-species vision-cone model. An agent 1 with orientation $\mathbf{\hat{u}_{1}}$ sees and interacts with other agents within a vision cone of full opening angle $2\vartheta$. Agent 2 is included in the perception zone of agent 1, such that there is a resulting nonreciprocal force acting on agent 1 which is directed along $\mathbf{r}=\mathbf{r}_{2}-\mathbf{r}_{1}$. There is also a torque turning its orientation to be aligned with the orientation $\mathbf{\hat{u}_{2}}$ of the second agent. In this example, agent 2 with its own vision cone does not see agent 1 and thus there is no interaction between them that influences the motion of agent 2. Likewise, agent 3 is not visible for agent 1 such that there is no mutual interaction between agents 1 and 3 that influences the motion of either of them.
II Active Model N
II.1 Microscopic basis
@@
We start from a microscopic picture of particles or agents interacting through vision cones, as illustrated in Fig. 1. Consider a general form of pairwise nonreciprocal force between single-component particles with an orientational degree of freedom on a plane, i.e.,
[…]
@@
[…]
@@
$f(r,\cos\varphi_{1},\cos\varphi_{2})=h(r)\Theta(-\cos\vartheta+\cos\varphi_{1}),$ (4)
@@
$M(r,\hat{\mathbf{u}}_{1},\hat{\mathbf{u}}_{2})=\tau(r)\Delta\varphi_{21}\Theta(-\cos\vartheta+\cos\varphi_{1}),$ (5)
@@
Our main focus here is on establishing the corresponding continuum field theory for single-species nonreciprocity, and predicting the emergence of new patterns and dynamics particularly those with structural chirality and parity-time (PT) symmetry breaking. To derive the continuum theory, we start from the microscopic Langevin equations containing the interparticle forces and torques. The dynamics of the full one-particle density field $\varrho(\mathbf{r},\hat{\mathbf{u}},t)$ is then derived by integrating out the stochastically equivalent Smoluchowski equation. By a systematic expansion in terms of a particle density field $\rho(\mathbf{r},t)$ and a polarization density field $\mathbf{P}(\mathbf{r},t)$, the coarse-grained continuum field equations can be obtained.
II.2 Derivation of generic continuum field theory
In Appendix A we construct a general nonreciprocal continuum field theory based on this approach, before applying the specific forms of nonreciprocal forces and torques given above to derive Active Model N. In the following we explain the main idea of the derivation. Similar strategies have been successfully applied to other physical systems, including the different case for self-propelled particles [27].
We consider a system of two-dimensional (2D) Brownian particles. The position $\mathbf{r}_{i}$ and orientation $\phi_{i}$ of the $i$th particle obey the Langevin equations
$\dot{\mathbf{r}}_{i}(t)$ $=\beta D_{\mathrm{T}}\mathbf{F}(\{\mathbf{r}_{j}(t),\phi_{j}(t)\})+\sqrt{2D_{\mathrm{T}}}\bm{\xi}_{i}(t),$ (6)
$\dot{\phi}_{i}(t)$ $=\beta D_{\mathrm{R}}M(\{\mathbf{r}_{j}(t),\phi_{j}(t)\})+\sqrt{2D_{\mathrm{R}}}\chi_{i}(t),$ (7)
where $\beta=1/(k_{\mathrm{B}}T)$ with Boltzmann constant $k_{\mathrm{B}}$ and temperature $T$, $D_{\mathrm{T}}$ and $D_{\mathrm{R}}$ represent the translational and rotational diffusion coefficients, $\mathbf{F}$ is the interaction force, $M$ is the interaction torque, and $\bm{\xi}_{i}$ and $\chi_{i}$ are white noises with zero mean and unit variance. Here the dot denotes a time derivative. This microscopic model is generic, since no restrictions have been made on force and torque apart from the assumption that they are pairwise (without three- or higher-body interactions). We deliberately do not include any self-propulsion terms to make sure that any nonequilibrium effects observed are just a consequence of nonreciprocal interactions. This allows us to study and understand the effect of nonreciprocity in isolation and to build a continuum model that incorporates single-species nonreciprocity in a minimal way (similar to the nonreciprocal Cahn-Hilliard equation [19, 22, 25, 39] which constitutes a minimal model for the case of multispecies nonreciprocity).
We then derive the corresponding Smoluchowski equation which is stochastically equivalent to Eqs. (6) and (7), integrate over the degrees of freedom of all particles except for one. This gives the dynamic equation for the orientation-dependent one-body density $\varrho$, i.e.,
$\frac{\partial}{\partial t}\varrho(\mathbf{r},\hat{\mathbf{u}},t)=D_{\mathrm{T}}\bm{\nabla}^{2}\varrho(\mathbf{r},\hat{\mathbf{u}},t)+D_{\mathrm{R}}\partial_{\phi}^{2}\varrho(\mathbf{r},\hat{\mathbf{u}},t)$
$\qquad+\beta D_{\mathrm{T}}\bm{\nabla}\cdot\bigg\{\varrho(\mathbf{r},\hat{\mathbf{u}},t){\int\;\;\>\hskip-5.83331pt\mathrm{d}^{2}r^{\prime}}\,{\int\;\;\>\hskip-5.83331pt\mathrm{d}^{2}u^{\prime}}\,$
$\qquad\Big[-\mathbf{F}(\mathbf{r},\mathbf{r}^{\prime},\hat{\mathbf{u}},\hat{\mathbf{u}}^{\prime},t)g(\mathbf{r},\mathbf{r}^{\prime},\hat{\mathbf{u}},\hat{\mathbf{u}}^{\prime},t)\varrho(\mathbf{r}^{\prime},\hat{\mathbf{u}}^{\prime},t)\Big]\bigg\}$
$\qquad+\beta D_{\mathrm{R}}\partial_{\phi}\bigg\{\varrho(\mathbf{r},\hat{\mathbf{u}},t){\int\;\;\>\hskip-5.83331pt\mathrm{d}^{2}r^{\prime}}\,{\int\;\;\>\hskip-5.83331pt\mathrm{d}^{2}u^{\prime}}\,$
$\qquad\Big[-M(\mathbf{r},\mathbf{r}^{\prime},\hat{\mathbf{u}},\hat{\mathbf{u}}^{\prime},t)g(\mathbf{r},\mathbf{r}^{\prime},\hat{\mathbf{u}},\hat{\mathbf{u}}^{\prime},t)\varrho(\mathbf{r}^{\prime},\hat{\mathbf{u}}^{\prime},t)\Big]\bigg\},$ (8)
where $\mathbf{F}$ and $M$ are the force and torque that a particle at position $\mathbf{r}^{\prime}$ with orientation $\hat{\mathbf{u}}^{\prime}$ exerts on a particle at position $\mathbf{r}$ with orientation $\hat{\mathbf{u}}$, and $g$ is the pair-distribution function [40], which measures how the probability of finding a particle at position $\mathbf{r}^{\prime}$ with orientation $\hat{\mathbf{u}}^{\prime}$ is modified by the presence of a particle at position $\mathbf{r}$ with orientation $\hat{\mathbf{u}}$.
A general difficulty in deriving field theories for interacting particles is the requirement of an expression for the pair-distribution function $g$. This problem can be solved in the passive case by approximating it via relations known from equilibrium density functional theory [41] and in the active case by exploiting the fact that $g$ is known from simulation results [42, 43]. However, neither of these is possible here since the system is considerably out of equilibrium and since $g$ has never been studied for systems of this type governed by nonreciprocity. Hence, we use the mean-field approximation, by simply setting $g=1$. This is better justified for soft interaction potentials (as used in this study; see Sec. II.3), for which in the passive case the mean-field approximation is known to be very accurate [44]. Equation (8) is then reduced to
$\frac{\partial}{\partial t}\varrho(\mathbf{r},\hat{\mathbf{u}},t)=\,$ $D_{\mathrm{T}}\bm{\nabla}^{2}\varrho(\mathbf{r},\hat{\mathbf{u}},t)+D_{\mathrm{R}}\partial_{\phi}^{2}\varrho(\mathbf{r},\hat{\mathbf{u}},t)$
$\!+\beta D_{\mathrm{T}}\bm{\nabla}\cdot\bigg\{\varrho(\mathbf{r},\hat{\mathbf{u}},t){\int\;\;\>\hskip-5.83331pt\mathrm{d}^{2}r^{\prime}}\,{\int\;\;\>\hskip-5.83331pt\mathrm{d}^{2}u^{\prime}}\,$
$\quad\Big[-\mathbf{F}(\mathbf{r},\mathbf{r}^{\prime},\hat{\mathbf{u}},\hat{\mathbf{u}}^{\prime},t)\varrho(\mathbf{r}^{\prime},\hat{\mathbf{u}}^{\prime},t)\Big]\bigg\}$
$\!+\beta D_{\mathrm{R}}\partial_{\phi}\bigg\{\varrho(\mathbf{r},\hat{\mathbf{u}},t){\int\;\;\>\hskip-5.83331pt\mathrm{d}^{2}r^{\prime}}\,{\int\;\;\>\hskip-5.83331pt\mathrm{d}^{2}u^{\prime}}\,$
$\quad\Big[-M(\mathbf{r},\mathbf{r}^{\prime},\hat{\mathbf{u}},\hat{\mathbf{u}}^{\prime},t)\varrho(\mathbf{r}^{\prime},\hat{\mathbf{u}}^{\prime},t)\Big]\bigg\}.$ (9)
Note that Eq. 8 actually coincides with Eq. 9 if we define a modified interaction force $\tilde{\mathbf{F}}=g\mathbf{F}$ and a modified torque $\tilde{M}=gM$. Thus, a derivation without the mean-field approximation would essentially correspond to using the modified form of interactions, and since our results do not depend on specific details of the interactions, they are expected to be fairly general. (This would not be the case for, e.g., motility-induced phase separation in active spheres, since there $g$ has a different angular dependence than $\mathbf{F}$ and using $\tilde{\mathbf{F}}$ rather than $\mathbf{F}$ changes the underlying symmetries. In the setup here, $\mathbf{F}$ and $M$ already have the most general angular dependence compatible with the symmetries of the system, implying that our result will also be general.)
Given global translational and rotational invariance, $\mathbf{F}$ and $M$ can be written as $\mathbf{F}=f(r,\varphi_{1},\varphi_{2})\mathbf{\hat{u}}_{\mathrm{F}}$ and $M=M(r,\varphi_{1},\varphi_{2})$ with orientation angles $\varphi_{1}=\phi_{\mathrm{R}}-\phi$, $\varphi_{2}=\phi^{\prime}-\phi$, and $\mathbf{r}^{\prime}-\mathbf{r}=r\mathbf{\hat{u}}(\phi_{\mathrm{R}})$. The scalar functions $f$ and $M$ depend only on rotationally invariant quantities, whereas the direction of $\mathbf{F}$, given by the unit vector $\mathbf{\hat{u}}_{\mathrm{F}}$, is obviously not rotationally invariant. We now assume $\mathbf{\hat{u}}_{\mathrm{F}}=\mathbf{\hat{u}}(\phi_{\mathrm{R}})$ (cf. Eq. 1), as is the case for many pairwise interaction forces usually considered.
The rest of the derivation is rather lengthy and discussed in detail in Appendix A. Here, we list the key steps of the derivation as an overview: (i) A truncated Fourier expansion is performed for the dependence of $f$ and $M$ on angles $\varphi_{1}$ and $\varphi_{2}$. (ii) A Cartesian orientational expansion [45, 46] is performed for the dependence of $\varrho$ on $\mathbf{\hat{u}}$, which allows for the replacement
[…]
where “:” is a double tensor contraction. (iii) A truncated gradient expansion [47, 27] is performed for the spatially nonlocal terms in Eq. 9, which allows to express them as a sum of local terms. (iv) We evaluate the angular integrals, which can be done analytically in closed form, to obtain coupled dynamical equations for the fields $\rho$, $\mathbf{P}$, and $\underline{\mathbf{Q}}$. (v) We then make a quasi-stationary approximation [27] $\partial\underline{\mathbf{Q}}/\partial t=\underline{\mathbf{0}}$, solve the resulting equation for $\underline{\mathbf{Q}}$, and insert the result into the dynamical equations for $\rho$ and $\mathbf{P}$ to obtain closed equations of motion for these two fields.
This results in a general field-theoretical model (given by Eqs. (60) and (61) in Appendix A) which is valid for any reciprocal or nonreciprocal interaction forces and torques. Note that the derivation is predictive, i.e., it provides explicit microscopic expressions for all model coefficients.
II.3 Nonreciprocal field theory: Active Model N
We can derive Active Model N from the general theoretical model (Eqs. (60) and (61)) simply by explicitly calculating the model coefficients using the nonreciprocal force and torque equations (4) and (5). For this choice of force and torque, most of the coefficients in the general model are equal to zero, such that the theory considerably simplifies. The model is further simplified by dropping the terms of second order contribution (see Appendix A). Finally, the resulting model is nondimensionalized via a length scale $a$ representing the interaction range of force and a time scale $a^{2}/D_{\mathrm{T}}$, with density fields $\rho$, $\mathbf{P}$, and $\underline{\mathbf{Q}}$ rescaled by $a^{2}$. We then reach a new continuum density-field theory, i.e., Active Model N,